OpenStudy (anonymous):

suppose a triangle has two sides of length 2 and 3 and that the angle between these two sides is 60 degrees. What is the length of the third side of the triangle?

OpenStudy (welshfella):



OpenStudy (welshfella):

use the Cosine Rule c^2 = a^2 + b^2 - 2ab cos C here we have c = x , a = 2 b = 3 and C = 60 degrees